# Wenfan Ou and Sheng Bi

### v3

@@ 5 Discrepancy-Based RMDPs
Besides estimating the moment information from historical observations, the decision-maker can also obtain a nominal/reference distribution to approximate the underlying probability distribution, either directly from historical data or expert experience. If data or expert experience is considerably reliable, e.g., high-quality data and recognizable patterns of uncertainty, it is reasonable to believe that the discrepancy between the nominal distribution and the true distribution is sufficiently small. More precisely, such an ambiguity set can be depicted in the following generic form:
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}\left(\Xi,\mathcal{F}\right)\left|~{}D(\hat{P},P)\leq\theta\right\}\right.,$ (14)
where $\hat{P}$ denotes the nominal probability measure, and $D(\cdot,\cdot):\mathfrak{M}\left(\Xi,\mathcal{F}\right)\times\mathfrak{M}\left(\Xi,\mathcal{F}\right)\rightarrow\mathbb{R}_{+}\cup\{\infty\}$ is a function that measures the discrepancy between two probability measures $P,~{}\hat{P}\in\mathfrak{M}\left(\Xi,\mathcal{F}\right)$. The parameter $\theta\in[0,\infty]$, also called the level of robustness, controls the size of the ambiguity set. Same as before, we abuse the notation $\mathcal{P}$ to denote the ambiguity set of corresponding $\mathbf{p}$ or $\mu$ with respect to the probability measure $P$. In line with Rahimian & Mehrotra (2019), we refer to RMDPs with ambiguity sets like (14) as discrepancy-based RMDPs, and present several prevalent ambiguity sets in this section.
@@ 5.1 Norm Ambiguity
Norm ambiguity sets, typically constructed on a probability simplex, measure the discrepancy between two distributions using various norms. While this construction appears analogous to (Satia & Lave Jr 1973), it emphasizes the difference between the entire distributions rather than individual probabilities. Generally, norm ambiguity sets can be represented as follows.
@@
For $s\in\mathcal{S},a\in\mathcal{A}$, the $s,a$-rectangular norm ambiguity sets can be defined as
$\mathcal{P}_{sa}\triangleq\left.\left\{\mathbf{p}\in\Delta^{|\mathcal{S}|}\right|\|\hat{\mathbf{p}}_{sa}-\mathbf{p}\|\leq\theta_{sa},~{}\theta_{sa}\geq 0\right\}.$ (15)
For $s\in\mathcal{S}$, the $s$-rectangular norm ambiguity sets can be defined as
$\mathcal{P}_{s}\triangleq\left\{\mathbf{p}_{s}=(\mathbf{p}_{s1},...,\mathbf{p}_{s|\mathcal{A}|})\in\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\left|\begin{array}[]{l}\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|},\quad\forall a\in\mathcal{A},\\ \textstyle\sum\limits_{a\in\mathcal{A}}\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|\leq\theta_{s},\theta_{s}\geq 0\end{array}\right\}\right.,$ (16)
@@
RMDPs with norm ambiguity sets can be easily formulated as convex programs, which makes them a preferred choice for large-scale problems. Among the various norms, the $L_{1}$-norm and $L_{\infty}$-norm have garnered significant attention in the literature.
@@
The $L_{1}$-norm (also referred to as variation distance) ambiguity sets are the most common and earliest studied norm in norm-constrained RMDPs literature (Iyengar 2005, Petrik & Subramanian 2014). The benefits of constructing such ambiguity sets are two-fold. Firstly, this construction enables the calculation of worst-case transition probabilities to be computed via linear programs, which brings computational efficiency. Secondly, the size of ambiguity sets can be easily devised using Hoeffding-style bounds with samples in the following proposition.
Proposition 5.2 ($L_{1}$-Norm Deviation Bound (Weissman et al. 2003))
@@
Let $n_{s,a}$ be the number of transitions from state $s$ by taking action $a$, and $\delta\in(0,1]$ be the confidence level. When the deviation bound $\theta_{s,a}$ is chosen as $\theta_{s,a}=\sqrt{\frac{2}{n_{s,a}}\log\frac{|\mathcal{S}||\mathcal{A}|2^{|\mathcal{S}|}}{\delta}}$, the underlying $P_{s,a}^{0}$ is contained in the ambiguity set with probability $1-\delta$.
Ho et al. (2018) consider $L_{1}$-norm constrained RMDPs with $(s,a)$-rectangular and $s$-rectangular assumption, respectively. For $(s,a)$-rectangular assumption, the authors propose a homotopy method, assuming that the ambiguity set emerges from the intersection of the probability simplex and a $w$-weighted $L_{1}$-norm ball, where the norm is defined as $\|x\|_{1,w}=\textstyle\sum_{i=1}^{n}w_{i}|x_{i}|$. The basic idea of homotopy methods is to trace the optimal solution to an optimization problem while increasing the value of some parameters. For $s$-rectangular assumption, the authors employ the bisection method, which has been explored in $(s,a)$-rectangular case (Nilim & El Ghaoui 2005), to solve robust Bellman update efficiently. Different from the previous work, they show there is an equivalent reformulation of the robust Bellman update, and the reformulation is easily computed via the bisection method.
While the homotopy method in $(s,a)$-rectangular case in (Ho et al. 2018) accelerates the search for the optimal solution, a linear program has to be solved for each state and each step of the value or policy iteration, which is costly for a large state space. To overcome this disadvantage, Ho et al. (2021) generalize the modified robust policy iteration in Kaufman & Schaefer (2013), proposing a partial policy iteration framework for $(s,a)$-rectangular ambiguity sets.
$L_{\infty}$-Norm is another approach to construct norm ambiguity sets, which aims to restrict the maximal deviation for each transition probability mass. Although $L_{\infty}$-norm is more natural and interpretable, there exist some challenges to employing $L_{\infty}$-norm compared to $L_{1}$-norm. Similar to Ho et al. (2018, 2021), Behzadian et al. (2021) also employ homotopy and bisection methods to design quasi-linear time complexity algorithms for $L_{\infty}$-constrained $(s,a)$-rectangular and $s$-rectangular RMDPs, respectively. However, previous results and guarantees established on the sparsity properties of $L_{1}$-norm do not hold for $L_{\infty}$-norm, leading to more complex algorithms.
@@ 5.2 $\phi$-Divergence Ambiguity
While norm ambiguity sets offer advantages in computational efficiency and modeling convenience, they are limited by their ability to provide probabilistic information and guarantees. As a result, there is growing interest in ambiguity sets designed with probabilistic metrics. $\phi$-divergence is one of the most popular class probabilistic metrics with tractable and desirably statistical properties. Here, we define general $\phi$-divergence ambiguity sets formally.
@@ A $\phi$-divergence ambiguity set is defined via
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}(\Xi,\mathcal{F})\left|~{}D^{\phi}(P,\hat{P})\leq\theta\right\}\right.,$ (17)
where $D^{\phi}(P,\hat{P}):=\int_{\Xi}\phi(\frac{\mathrm{d}P}{\mathrm{d}\hat{P}})\mathrm{d}\hat{P}$ is the similarity measure function whose concrete representation depends on divergence function ϕ():++{+}:italic-$\phi(\cdot):\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{+\infty\}$. The divergence function $\phi(\cdot)$ is convex, and satisfies $\phi(1)=0$, $00\phi(\frac{0}{0}):=$ 0 and $0\phi(\frac{a}{0}):=a\lim_{t\rightarrow\infty}\frac{\phi(t)}{t}$ for $a>0$ (Bayraksan & Love 2015, Pardo 2018).
Despite numerous $\phi$-divergences, only some divergences are widely accepted due to well-defined mathematical properties for robust optimization. Table 1 presents a list of typical divergence functions in probability style, where the second column $\phi(\cdot)$ is the corresponding $\phi$-divergence function, the third column $D^{\phi}(\mathbf{p}_{sa},\mathbf{q}_{sa})$ is the representation of $\phi$-divergence in probabilities style, and last column is the conjugate function of $\phi(\cdot)$.
@@
Kullback-Leibler (KL)-divergence is the most prevalent $\phi$-divergence in both DRO and RMDPs. As the cornerstone of modern RMDPs, Iyengar (2005) and Nilim & El Ghaoui (2005) both explore KL-divergence ambiguity sets to depict statistical uncertainty. By employing the Lagrangian duality, they reformulate the inner maximize problem in dual form,
@@
where $\lambda$ is the Lagrangian multiplier and $\hat{p}_{sas^{\prime}}$ is the nominal probability of transition $(s,a,s^{\prime})$. As the problem is reduced to a one-dimensional convex program, Nilim & El Ghaoui (2005) develop a bisection algorithm to obtain an $\epsilon$-optimal policy solved in polynomial time. To contribute to the marginally theoretical results, Iyengar (2005) proposes a conservative approximation ambiguity set in probability measure style,
@@
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}(\mathcal{S})\left|\sum\limits_{s\in\mathcal{S}}\frac{(P(s)-\hat{P}(s))^{2}}{\hat{P}(s)}\leq\theta\right\}\right.,$ (19)
and show that this approximation allows us to solve the inner optimization problem more efficiently.
Caro & Das Gupta (2022) incorporate findings from Iyengar (2005) and Nilim & El Ghaoui (2005) into robust multi-armed bandit (RMAB) problem. With KL-divergence ambiguity sets, they formulate the robust Bellman equation and introduce a heuristic policy to circumvent evaluating indices, which is a well-known policy in MAB but costly for RMAB. To counteract the curse of dimensionality, they propose an approximation method. This method begins with simulating the policy based on the nominal probabilities’ point estimates and defines the states that are not encountered during the simulation as low-priority. The expected reward of low-priority states is truncated to the immediate reward. Thereby, robust value iteration is applied exclusively to the high-priority states, significantly reducing computational costs.
@@
Similarly inspired by Iyengar (2005) and Nilim & El Ghaoui (2005), Liu et al. (2022) design a distributionally robust Q-learning algorithm with the KL divergence ambiguity sets. In virtue of the strong duality lemma from distributionally robust optimization under KL-perturbation (Hu & Hong 2013), the authors derive the distributionally robust Q-function, which transforms the infinite-dimensional problem into a finite-dimensional dual representation [^2] We display the distributionally robust Q-function without reward uncertainty and slightly modified from the original version of the paper.:
@@
$Q(s,a):=\sup_{\alpha,\beta\geq 0}\left\{R(s,a,s^{\prime})-\alpha\theta-\gamma\beta\log\left(\mathbb{E}_{\hat{\mathbf{p}}_{s,a}}\left[\exp\left(-\frac{\max_{b\in\mathcal{A}}Q(s^{\prime},b)}{\beta}\right)\right]\right)-\gamma\beta\theta\right\}.$ (20)
The authors claim that the plug-in estimator of the distributionally robust Bellman Operator is biased because of the non-linearity. Hence, they leverage the multi-level Monte-Carlo scheme (Blanchet & Glynn 2015) to estimate the unbiased dual value using samples from the simulator and show that the proposed algorithm converges asymptotically to the optimal distributionally robust problem.
@@
$\chi^{2}$-divergence, also known as $\chi^{2}$-test, is also commonly used to model the ambiguity sets in RMDPs. Compared to KL-divergence, $\chi^{2}$- divergence possesses several advantageous properties, including symmetry and reduced sensitivity to minor differences. Hanasusanto & Kuhn (2013) consider a data-driven stochastic control problem with continuous state and action spaces, where the transition probability matrix is estimated via Nadaraya-Watson (NW) kernel regression, as the nominal distribution. For instance, given $N$ observation histories $\{\xi^{i}_{t}\}^{T}_{t=1},i\in[N]$, the conditional expectation can be estimated as
@@
where $p_{ti}({\xi}_{t})=\frac{\mathrm{K_{NW}(\xi_{t}-\xi^{i}_{t})}}{\sum_{j=1}^{N}\mathrm{K_{NW}}(\xi_{t}-\xi^{j}_{t})},~{}i\in[N],t\in[T]$ and $\mathrm{K_{NW}}(\cdot)$ is the NW kernel density function. To mitigate the poor out-of-sample performance of the data-driven dynamic programming approach under sparse training data, the authors construct $\chi^{2}$-distance ambiguity sets. By general dual theorem and piece-wise linear or convex quadratic assumption of value-to-go function, they reduce the original problem to a tractable second-order cone program which can be solved efficiently with interior point algorithms. Instead of finding optimal policies, they design a robust data-driven dynamic program to approximate the value-to-go functions via interpolation.

### v4

@@ 5 Discrepancy-Based RMDPs
In addition to estimating (partial) moment information, it is possible for the decision-maker to obtain a nominal/reference distribution to approximate the underlying probability distribution, either directly from historical data or domain knowledge. If the available data or domain knowledge is reasonably reliable, such as high-quality data and recognizable patterns of uncertainty, it is reasonable to believe that the discrepancy between the nominal distribution and the true distribution is sufficiently small. Such a discrepancy-based ambiguity set can be represented by the following generic form:
$\mathcal{P}\triangleq\left\{\mathbb{P}\in\mathfrak{M}\left(\Xi,\mathcal{F}\right)\left|~{}D(\hat{\mathbb{P}},\mathbb{P})\leq\theta\right\}\right.,$ (23)
where $\hat{\mathbb{P}}$ denotes the nominal probability measure, and $D(\cdot,\cdot):\mathfrak{M}\left(\Xi,\mathcal{F}\right)\times\mathfrak{M}\left(\Xi,\mathcal{F}\right)\rightarrow\mathbb{R}_{+}\cup\{\infty\}$ is a function that measures the discrepancy between two probability measures ,^$\mathbb{P},~{}\hat{\mathbb{P}}\in\mathfrak{M}\left(\Xi,\mathcal{F}\right)$. The parameter $\theta\in[0,\infty]$, also called the level of robustness, limits the maximum discrepancy, thereby controlling the size of the ambiguity set. As before, we sightly abuse the notation $\mathcal{P}$ to denote the ambiguity set of corresponding $\mathbf{p}$ or $\mu$ with respect to the probability measure $\mathbb{P}$ for simplicity. In line with Rahimian & Mehrotra (2019), we refer to RMDPs with ambiguity sets like (23) as discrepancy-based RMDPs, and focus on several prevalent ambiguity sets in this section.
@@ 5.1 Norm Ambiguity
Norm ambiguity sets, typically constructed on a probability simplex, measure the discrepancy between two distributions regarding various norms. While this construction appears analogous to Satia & Lave Jr (1973), it emphasizes the difference between the entire distributions rather than individual probabilities. Generally, norm ambiguity sets can be represented as follows.
@@
$\mathcal{P}_{sa}\triangleq\left.\left\{\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}\right|\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|\leq\theta_{sa},~{}\theta_{sa}\geq 0\right\},$ (24)
where $\hat{\mathbf{p}}_{sa}$ is the nominal distribution over the next state concerning state-action pair $(s,a)$, and $\mathbf{p}_{sa}$ belongs to $|\mathcal{S}|$-dimension probability simplex. The deviation with respect to the predetermined norm is constrained to be no larger than a parameter $\theta_{sa}$. For $s\in\mathcal{S}$, a $s$-rectangular norm ambiguity set can be defined similarly as
$\mathcal{P}_{s}\triangleq\left\{\mathbf{p}_{s}=(\mathbf{p}_{s1},...,\mathbf{p}_{s|\mathcal{A}|})\in\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\left|\begin{array}[]{l}\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|},\ \forall a\in\mathcal{A},\\ \textstyle\sum\limits_{a\in\mathcal{A}}\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|\leq\theta_{s},\ \theta_{s}\geq 0\end{array}\right\}\right.,$ (25)
where $\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|},\ \forall a\in\mathcal{A}$ ensures that $\mathbf{p}_{sa}$ is a valid probability distribution over the next state, while $\sum_{a\in\mathcal{A}}\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|\leq\theta_{s}$ establishes that the sum of the distances between the nominal probabilities $\hat{\mathbf{p}}_{sa}$ and plausible probabilities $\mathbf{p}_{sa}$ across all actions is bounded by $\theta_{s}$.
@@
Common norms, such as the $L_{1}$, $L_{\infty}$, and $L_{p}$-norm ($1<p<\infty$), are known to be convex. The convexity allows RMDPs with norm-based ambiguity sets to be readily formulated as convex programs, making them a preferred choice for tackling large-scale problems. Moreover, these norms often have intuitive geometric interpretations, which assist the decision-maker in establishing a desirable model. Among the various norms, the $L_{1}$ and $L_{\infty}$ norms have received significant attention in the literature.
@@
The $L_{1}$-norm, also known as the variation distance, is the most widely used and one of the earliest studied norms in the context of norm-constrained RMDPs (Iyengar 2005, Petrik & Subramanian 2014). The construction of $L_{1}$-norm-based ambiguity sets offers two significant advantages. First, this construction enables the calculation of worst-case transition probabilities to be computed via linear programs, which brings computational efficiency. Second, the size of these ambiguity sets can be determined using Hoeffding-style bounds, which limit the probability of large deviations in the sums of random variables. This property is formalized in the following proposition.
Proposition 5.2 ($L_{1}$-Norm Finite Sample Bound (Wiesemann et al., 2003))
@@
Let $n_{sa}$ be the number of transitions from state $s$ by taking action $a$ and $\delta\in(0,1]$ be the confidence level. When the deviation bound parameter $\theta_{sa}$ is chosen as $\theta_{sa}=\sqrt{\frac{2}{n_{sa}}\log\frac{|\mathcal{S}||\mathcal{A}|2^{|\mathcal{S}|}}{\delta}}$, the underlying $\mathbf{p}_{sa}^{0}$ is contained in the ambiguity set with probability $1-\delta$.
Note that the bound in Proposition 5.2 gets tighter as $n_{sa}$ increases, reflecting higher confidence with more observations. The bound also depends on the size of the state and action spaces, as well as the desired confidence level $\delta$, all of which contribute to the complexity of achieving an accurate approximation. This bound allows for the data-driven construction of ambiguity sets by providing a principled way to balance between robustness and conservatism based on the amount of available data.
Ho et al. (2018) consider a $w$-weighted $L_{1}$-norm constrained RMDPs with $(s,a)$- and $s$-rectangular assumption, where the norm is defined as $\|x\|_{1,w}=\textstyle\sum_{i=1}^{n}w_{i}|x_{i}|$. In this paper, the authors leverage the properties of $L_{1}$-norm to develop fast Bellman updates, rather than directly using LPs. Specifically, the authors start from the Q-function $q\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ of RMDPs where:
$q_{sa}(\eta)=\min_{\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}}\{R(s,a)+\gamma\cdot\mathbf{p}_{sa}^{T}v:\|\mathbf{p}_{sa}-\hat{\mathbf{p}}_{sa}\|\leq\eta\}$ (26)
Where $R(s,a)$ is an immediate reward unrelated to the next state, $v:=(V(1),...,V(|\mathcal{S}|))$ denotes the state value vector and $\eta$ is controlled parameter and is smaller than predetermined robustness budget $\theta_{sa}$. As the robust Bellman optimality equation holds under rectangular assumptions (Iyengar 2005), the authors derive optimal Q-function $q^{*}$ in return. Benefiting from the structure of $L_{1}$-norm, problem (26) can be readily reformulated as a linear program:
$\begin{array}[]{lll}q_{sa}(\eta)=&\min\limits_{\mathbf{p}_{sa}\in\mathbb{R}^{|\mathcal{S}|},~{}l\in\mathbb{R}^{|\mathcal{S}|}}&\mathbf{p}_{sa}^{T}(R(s,a)\mathbf{1}+\gamma v)\\ &\text{s.t.}&\mathbf{p}_{sa}-\hat{\mathbf{p}}_{sa}\leq l\\ &&\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\leq l\\ &&\mathbf{p}_{sa}\geq\mathbf{0}\\ &&\mathbf{1}^{\mathsf{T}}\mathbf{p}_{sa}=1,\quad w^{\mathsf{T}}l=\eta\end{array}$ (27)
Given the LP in (27) and $(s,a)$-rectangular assumption, the authors develop a homotopy method which starts from a trivial feasible solution $\eta=0$ (i.e., $\mathbf{p}_{sa}=\hat{\mathbf{p}}_{sa}$) and then track the optimal solution $\mathbf{p}_{sa}$ as $\eta$ gradually increases. Since $q_{sa}(\eta)$ and $\mathbf{p}_{sa}$ are both piecewise linear in $\eta$, the proposed homotopy approach can be efficiently implemented by tracing the basic feasible solutions. For the $s$-rectangular case, the authors show that an equivalent reformulation of the robust Bellman update exists, as follows:
$V^{*}(s)=\max_{\pi\in\Delta^{|\mathcal{S}|}}\min_{\eta\in\mathbb{R}^{|\mathcal{A}|}}\left\{\sum_{a\in\mathcal{A}}\pi(s,a)q_{sa}(\eta_{a}):\sum_{a\in\mathcal{A}}\eta_{a}\leq\theta_{s}\right\}\iff\min_{u_{s}\in\mathbb{R}}\left\{u_{s}:\sum_{a}q_{sa}^{-1}(u_{s})\leq\theta_{s}\right\}$ (28)
where $q_{sa}^{-1}(u)=\min_{\mathbf{p}_{sa}\in\Delta^{S}}\left\{\|\mathbf{p}_{sa}-\hat{\mathbf{p}}_{sa}\|_{1,w_{a}}:r_{a}+\gamma\mathbf{p}_{sa}^{\mathsf{T}}v\leq u_{s}\right\}$. This reformulation is non-trivial; therefore, we provide the intuition behind (28) instead of a concrete proof here. For a given state $s$, with a limited robustness budget $\theta_{s}$, $q_{sa}^{-1}(u)$ can be interpreted as the minimum robustness budget assigned to action $a$ such that the value-to-go function does not exceed $u_{s}$. Thus, minimizing $u_{s}$ determines the worst-case transition probabilities $\mathbf{p}_{sa}$ that lead to the lowest value-to-go function, which is equivalent to optimizing the robust Bellman equation. Note that this reformulation simplifies the original update into a one-dimensional problem, and the authors propose a bisection algorithm atop the homotopy method to efficiently solve it.
Although the homotopy method for the $(s,a)$-rectangular case in Ho et al. (2018) has accelerated the search for the optimal solution, a linear program has to be solved for each state and each step of the value or policy iteration, which is costly in large state space. To overcome this disadvantage, Ho et al. (2021) propose a partial policy iteration (PPI) framework for $(s,a)$- or $s$-rectangular ambiguity sets. In contrast to general robust policy iteration procedures (Iyengar 2005, Nilim & El Ghaoui 2005), the PPI simplifies the policy evaluation step where it only solves a regular ordinary MDP (which is constructed from the corresponding RMDP), thus any advanced MDP algorithm can be applied directly, making the iteration procedures more efficiently. Combining PPI with the homotopy method or bisection method, the resulting algorithms achieve a significant speedup. Additionally, the PPI proposed in this paper is the first policy iteration method that provably converges to the optimal solution for s-rectangular RMDPs, extending the previous findings that PPI merely holds for $(s,a)$-rectangular RMDPs (Kaufman & Schaefer 2013).
The $L_{\infty}$-Norm offers an alternative approach for constructing norm-based ambiguity sets, aiming to constrain the maximum deviation of each transition probability mass. Although $L_{\infty}$-norm is more intuitive and interpretable, and the corresponding ambiguity sets empirically outperform $L_{1}$-norm-based ambiguity sets in some circumstances (Behzadian et al., 2021), it presents certain challenges compared to the $L_{1}$-norm. First, the concentration inequalities that facilitate the data-driven construction of high-confidence RMDPs do not hold. Second, most efficient algorithms, such as those in Ho et al. (2018, 2021), rely on the sparsity properties of $L_{1}$-norm and therefore cannot be directly applied to $L_{\infty}$-norm ambiguity sets.
Inspired by Ho et al. (2018, 2021), Behzadian et al. (2021) also employ homotopy and bisection methods to design quasi-linear time complexity algorithms for $L_{\infty}$-constrained $(s,a)$ and $s$-rectangular RMDPs. The key procedure involves reformulating the Q-function as a parametric LP similar to (27), but replacing the bounded constraints with
$\mathbf{1}^{T}\mathbf{p}_{s}a=1,-\eta\leq p_{sas^{\prime}}-\hat{p}_{sas^{\prime}}\leq\eta,\ p_{sas}\geq 0,\ \forall s^{\prime}\in\mathcal{S}.$
Hence, leveraging the structural properties of the new LPs, such as the piecewise linear and non-increasing property of $q_{sa}(\eta)$ with respect to $\eta$, the authors modify the homotopy and bisection methods to accommodate the constraints brought by $L_{\infty}$-norm.
@@ 5.2 $\phi$-Divergence Ambiguity
While norm ambiguity sets offer advantages in computational efficiency and modeling convenience, they are limited by their ability to convey probabilistic information and guarantees. As a result, there is growing interest in ambiguity sets designed with probabilistic metrics. Among these, $\phi$-divergence stands out as a widely favored class due to its tractable and desirable statistical characteristics. Here, we formally define general $\phi$-divergence ambiguity sets.
@@ A $\phi$-divergence ambiguity set is defined via
$\mathcal{P}\triangleq\left\{\mathbb{P}\in\mathfrak{M}(\Xi,\mathcal{F})\left|~{}D^{\phi}(\mathbb{P},\hat{\mathbb{P}})\leq\theta\right\}\right.,$ (29)
where $D^{\phi}(\mathbb{P},\hat{\mathbb{P}}):=\int_{\Xi}\phi(\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\hat{\mathbb{P}}})\mathrm{d}\hat{\mathbb{P}}$ is the similarity measure function whose concrete representation depends on divergence function ϕ():++{+}:italic-$\phi(\cdot):~{}\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{+\infty\}$. The divergence function $\phi(\cdot)$ is required to be convex, and satisfies $\phi(1)=0$, $00\cdot\phi(\frac{0}{0}):=$ 0 and $0\cdot\phi(\frac{a}{0}):=a\lim_{t\rightarrow\infty}\frac{\phi(t)}{t}$ for $a>0$ (Bayraksan & Love 2015, Pardo 2018).
Despite numerous $\phi$-divergences, only some divergences are widely accepted due to their well-defined mathematical properties for robust optimization. Table 1 presents a list of typical divergence functions in terms of probabilities, where the second column $\phi(\cdot)$ is the corresponding $\phi$-divergence function, the third column $D^{\phi}(\mathbf{p}_{sa},\mathbf{q}_{sa})$ is the representation of $\phi$-divergence in terms of probabilities, and last column is the conjugate function of $\phi(\cdot)$.
@@
Kullback-Leibler (KL) divergence, known for its interpretability and convexity, is among the most widely used $\phi$-divergences in both DRO and RMDPs. As the cornerstone of modern RMDPs, Iyengar (2005) and Nilim & El Ghaoui (2005) both improve upon traditional MDPIPs by exploring KL-divergence ambiguity sets to depict statistical uncertainty better. While sharing the same support between estimates and true randomness, KL-divergence limits the deviation in terms of the entire transition probability distribution, rather than individual probability mass. With $(s,a)$-rectangular assumption, they reformulate the inner maximization problem in dual form by employing the standard Lagrangian duality, as follows:
@@
where $\lambda$ is the Lagrangian multiplier and $\hat{p}_{sas^{\prime}}$ is the nominal probability of transition $(s,a,s^{\prime})$. As the problem is reduced to a one-dimensional convex program with respect to $\lambda$, Nilim & El Ghaoui (2005) develop a bisection algorithm to obtain an $\epsilon$-optimal policy solved in polynomial time. Nonetheless, the reformulation problem still involves $\log$ and $\exp$ operators, posing certain computational difficulties. To refine the theoretical foundations, Iyengar (2005) utilizes the fact that $\log(1+x)\leq x$ for all $x\in\mathbb{R}$, and proposes a conservative approximation ambiguity set in terms of probability measure,
@@
$\mathcal{P}\triangleq\left\{P_{sa}\in\mathfrak{M}(\mathcal{S},\mathcal{F}^{\mathcal{S}})\ \middle|\ \sum\limits_{s^{\prime}\in\mathcal{S}}\frac{(P_{sa}(s^{\prime})-\hat{P}_{sa}(s^{\prime}))^{2}}{\hat{P}_{sa}(s^{\prime})}\leq\theta\right\},$ (31)
where $\mathcal{F}^{\mathcal{S}}$ is the $\sigma$-algebra associated with sample space $\mathcal{S}$ and the probability measure $P$ belongs to probability space $\mathfrak{M}(\mathcal{S},\mathcal{F}^{\mathcal{S}})$ that assigns probabilities to next state $s^{\prime}$ occurring within a single period transition $(s,a,s^{\prime})$ (with the time subscript omitted). Essentially, the summation in (31) represents a modified $\chi^{2}$-distance between $P$ and $\hat{P}$, the nominal probability measure. The author shows that this approximation allows us to exactly solve the inner optimization problem with the complexity $\mathcal{O}(|\mathcal{S}|\log|\mathcal{S}|)$, often more efficiently than original KL-divergence ambiguity set with complexity $\mathcal{O}(|\mathcal{S}|^{1.5}\log|V_{max}/\epsilon|)$ where the solution is $\epsilon$-optimal.
@@
Motivated by the findings in Iyengar (2005) and Nilim & El Ghaoui (2005), Liu et al. (2022) design a distributionally robust Q-learning algorithm with the KL-divergence ambiguity sets. Like (30), the authors leverage the existence of the robust Bellman equation (arising from $(s,a)$-rectangular assumption) and the strong duality lemma from classical DRO results under KL-perturbation (Hu & Hong 2013) to derive a distributionally robust Q-function, which transforms the primal infimum (inner) infinite-dimensional problem into a supremum finite-dimensional dual representation [^2] For simplicity, we demonstrate the distributionally robust Q-function using a simplified formulation without reward uncertainty that is slightly modified from the original version of the paper.:
@@
$Q(s,a):=R(s,a)+\gamma\sup_{\lambda\geq 0}\left\{-\lambda\log\left(\mathbb{E}_{\hat{\mathbf{p}}_{s,a}}\left[\exp\left(-\frac{\max_{b\in\mathcal{A}}Q(s^{\prime},b)}{\lambda}\right)\right]\right)-\lambda\theta\right\}.$ (32)
where $R(s,a)$ is the immediate reward unrelated to the next state and $\gamma$ is the discount factor. The supremum is taken over the non-negative dual variable $\lambda$ derived from the primal problem transformation. The key term $\lambda\log(\mathbb{E}[\cdot])$ results from the dual formulation of the distributionally robust problem, where the expectation taken over the nominal transition probabilities $\hat{\mathbf{p}}_{s,a}$, ensuring that the reformulation becomes finite-dimensional and tractable. In essence, the worst-case Q-function is equivalent to a plug-in estimator over nominal transition probabilities when the deviation measured by the KL-divergence is not larger than $\theta$.
[…]
@@
Another commonly used approach to model the ambiguity sets in RMDPs is $\chi^{2}$-divergence. Compared to KL-divergence, $\chi^{2}$-divergence possesses several advantageous properties, including symmetry, reduced sensitivity to minor differences, and computational convenience (recall the conservative approximation proposed by Iyengar (2005) above is essentially $\chi^{2}$-divergence). Hanasusanto & Kuhn (2013) consider a data-driven stochastic control problem with continuous state and action spaces, where the transition probability matrix is estimated via Nadaraya-Watson (NW) kernel regression, as the nominal distribution. Given finite horizon $T$ and $N$ sample trajectories $\{\xi^{i}_{t}\}^{T}_{t=1},~{}i\in[N]$, the authors use empirical estimates to approximate the conditional expected value of future state $V_{t+1}$ concerning current observation $\xi_{t}$, as follows:
@@
where $p_{ti}({\xi}_{t})=\frac{\mathrm{K_{NW}(\xi_{t}-\xi^{i}_{t})}}{\sum_{j=1}^{N}\mathrm{K_{NW}}(\xi_{t}-\xi^{j}_{t})},~{}i\in[N],t\in[T]$ and $\mathrm{K_{NW}}(\cdot)$ is the NW kernel density function. In the right-hand side of (33), the weights give more importance to samples that are “closer” to the current observation $\xi_{t}$ in the future space. This method provides a non-parametric way to estimate conditional expectations, which is particularly useful in continuous state spaces where traditional tabular methods might not be applicable. The NW kernel regression allows for smooth interpolation between observed data points, providing estimates even for states not directly observed in the training data.
@@
However, if the training data is sparse, the NW estimates typically exhibit high variance, leading to the poor out-of-sample performance of the data-driven DP. To mitigate the impact of this issue, the authors construct $\chi^{2}$-distance ambiguity sets. Implicitly assuming $(s,a)$-rectangular assumption, they reformulate the inner infinite-dimensional maximum problem as a tractable minimum problem by standard dual theory. By exploiting the structure of $\chi^{2}$-distance ambiguity sets, the reformulation can be expressed as a second-order cone program (SOCP) under specific conditions. For instance, if the value function is piecewise linear or convex quadratic, and other mild assumptions hold—such as the immediate reward or cost function being convex quadratic in the state and action—this SOCP can be efficiently solved using interior-point algorithms. Instead of finding the optimal policies, they design a robust data-driven dynamic program algorithm to approximate the value-to-go functions via interpolation.